Grade 12Trigonometry

Why tan(pie)x + tanx is non periodic....???? plz explain me!!

Abhijat  Pandey
15 Years agoGrade 12

1 Answer

Chetan Mandayam Nayakar
ApprovedApproved Tutor Answer15 Years ago

period of tan((pie)x) is 1 and that of tanx is π, the sum of two periodic functions is periodic, if an integral multiple of the period of the first function is equal to (either the same or different) integral multiple of the period of the second function.

so m*1=n*π,m/n=pi which is impossible. That is why it is non-periodic.